answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How do you find a b and c in a parabolic equation of y equals a times x to the power of 2 plus b times x plus c?

If you want the parameters a, b and c then you must know the values of x and y. Even in this case you will not be able to calculate unique values for a,b and c but rather get a set of answers which describes the relationship necessary between a,b and c that will satisfy your values for x and y.

Here is how to do that:

Start with the usual parabolic equation: y = ax^2 + bx + c

There is a standard formula to solve this equation for x when you want y to be zero (also called the 'roots' of the equation). That formula is x = (-b +/- sqrt ( 4ac - b^2 ) ) / 2a).

The sqrt means take the square root of what in parentheses. and +/- means the formula must be calculated twice, once using a plus in place of +/- and once using a minus.

Instead of y being zero we want to be able to make it any value we like but our formula only works for y being zero. So we need to subtract the y value from the original parabolic equation to make one side zero, that give us 0 = ax^2+bx + (c-y).

Now we can use the standard formula again , but this time with c-y in place of just c.

x = (-b +/- sqrt( 4a(c-y) - b^2 ) / 2a

Using this new formula, if you know x and y you can calculate a relationship between a,b and c that will always give you that value of x and y in a parabolic equation.

Why is the number 12 not matter but 12 marbles are matter?

because 12 is a number and marbles are things which is matter.

* * * * *


In much the same way as your thought about a marble is not matter - it is simply a concept in your mind. But a marble is matter.

What the z-score of a person with a raw score of 20?

There is not enough information to answer your question. To determine a Z-Score, the mean and standard deviation is also required.

What is the probability of getting a 2 3 or 12 on two dice?

Total number of possibilities on two dice = 36.

Number of ways to get 2, 3, or 12 = 4.

1, 1

1, 2

2, 1

6, 6

Probability = 4/36 = 1 chance in 9 = 0.111 = 11.1%

A number cube has 6 sides the sides are numberd 1 to 6if the cube is thrown oncewhat is the probability of rolling a odd number?

Chances=50% because on a 6 sided dice, as you have said, the sides are numbered 1,2,3,4,5 and 6. There are 3 even numbers and 3 odd numbers. Chances would be 3/6=50%

What does imprecation mean?

An imprecation is a curse. To imprecate is to call a curse upon someone or something.

What is the difference between probability distribution functions and probability density functions?

Probability density function (PDF) of a continuous random variable is a function

that describes the relative likelihood for this random variable to occur

at a point in the observation space.

The PDF is the derivative of the probability distribution (also known as

cummulative distriubution function (CDF)) which described the enitre range of values

(distrubition) a continuous random variable takes in a domain.

The CDF is used to determine the probability a continuous random variable occurs any (measurable) subset of that range.

This is performed by integrating the PDF over some range (i.e., taking the area under of CDF curve between two values).

NOTE: Over the entire domain the total area under the CDF curve is equal to 1.

NOTE: A continuous random variable can take on an infinite number of values. The probability that it will equal a specific value is always zero.

eg. Example of CDF of a normal distribution:

If test scores are normal distributed with mean 100 and standard deviation 10. The probability

a score is between 90 and 110 is:

P( 90 < X < 110 ) = P( X < 110 ) - P( X < 90 )

= 0.84 - 0.16 = 0.68.

ie. AProximately 68%.

How many 36's can you roll with two dice?

None.

Two ordinary dice can only sum to 12 and so you cannot roll a 36.

What are the odds of hitting 2 straits in a row at 5 card stud?

The odds of hitting 2 straights in a row at 5 card stud is about 5%.

Since the deck is reshuffled between deals, the two hands are sequentially unrelated, so the odds of two deals with two straights is simply th odds of one times the odds of another.

In order to draw a straight, you need to draw an ace, 2, 3, 4, 5, or a 10, jack, queen, king, ace, assuming that ace can be high or low.

The odds of drawing one of these cards is 10 in 13, but you have 5 chances to do so, so the odds of drawing the first card is 50 in 13.

The odds of drawing the second card is 39 in 51, but you have 4 chances to do so, so the odds of drawing the second card is 156 in 51.

Similarly, the odds of drawing the third, fourth, and fifith card is 38 in 50, 37 in 49, and 36 in 48.

Multiply these odds and you get 19,740,240 in 77,968,800. Reducing that to lowest terms, you get 246,753 in 1,112,110, which is the odds of drawing a straight. The numeric value of the odds is 0.22187823146990855221156180593646, or about 22%

Square that and you get 60,887,043,009 in 1,236,788,652,100. That is aready in lowest terms, so that is the odds of drawing two sequential straights. The numeric value of the odds is 0.049229949600214317813758989938262, or about 5%.

Is selecting 3 marbles at the same time independent or dependent event?

The second and third marbles depend upon the first, so they are dependent. Even if they are all being drawn together, because there is no replacement, each removal affects the original group of marbles.

How many was can 4 people sit in a row of 4 seats if two of them insist on sitting together?

they sit on top of eachover

and the 3rd guy gives the 1st guy a reach-around

and the 2nd guy gives the 3rd guy a blowjob

and the 4th guy videos it and puts it on youporn.

hope that helped.

How many ways can a combined 4 be rolled with two standard dice?

Assuming you mean a sum of 4, there are three ways this can be done: (1, 3), (2, 2) and (3, 1).

Five women live together in an apartment Two have blue eyes If two of the women are chosen at random what is the probability that they both have blue eyes?

well lets look at the probability on your first pick you will get a blue eyed girl

2 have blue out of 5, so yo uhave a 2/5 chance, or 40%

on your second pick, there are 4 girls left, and one of them will have blue eyes now

1 have blue out of 4, so 1/4 or 25%

so you multiply the two probabilities to get your final sequential probability

2/5 * 1/4

find a common denominattor, for this we will use 20

8/20 * 5/20 = 40/400

simplify it you get 1/10 0.10, or 10% chance

Trending Questions
Decision making theory? What are the probable sources of smoke? Probability of gettingm 2 tails when 2 coins are tossed? What is the state that exists when the probability of success is less than 100 percent? If you think you are a narcissist but you truly love others and feel suicidal at the thought of being one does this rule you out as one and can you train your mind into thinking another way? Are standards designed to evaluate price and quantity variances separately? What are the probabilities of being king? What is the chance of dying from pneumonia in 1920? How many ways can the letters apple be arranged? What is the probability that in a group of six people at least two were born in the same month? Taylor tossed one green number cube and one red number cube each cube is numbered from 1-6what is the probility that the sum of the numbers on the cube is 4 write the answer as a fraction in simplest? When you toss a coin three times the total number of possible outcomes is? Box A has 3 red marbles and 2 yellow marble Box B has 2 red marble and 1 yellow If you have to pick a red marble to win and can't look in the box which box would give you the best chance? The odds in favor of an event are 10 to1 ratio Find the probability that the event will occur? What are the chances of dying? How do you see that the sampling distribution of the mean is normal? Which shows the probability of choosing 2 blue socks from a drawer having 7 blue and 3 white socks? Eighty percent of all California drivers wear seat belts. if three drivers are pulled over what is the probability that all would be wearing their seat belts write as a percent to the nearest ten? Seven coins are tossed What is the probability of 3 heads or 2 tails? How does a bucket holds fifty marbles of red yellow and blue There are thirty red marbles and the probability of picking a yellow marble is three out of ten How many blue marbles are in the bucket?